OpenStudy (anonymous):

a Straight line with a slope of 3/8 passes through the point (-4, 0). use the definition of slope to find another point on the same line ( Hint: try graphing ) stilling dont understand

OpenStudy (mathstudent55):

Slope is the rise over the run = rise/run. Rise is how much you move along y (vertically) and run is how much you move along x (horizontally) from one point to another. A slope of 3/8 is the same as a slope of (-3)/(-8) since (-3)/(-8) = 3/8 See graph below: